OpenStudy (anonymous):

For which of the following functions does f(a)+f(b) = f(a+b)? A: y=2x+3 B: y=2x C: y=x^2 D: y=sqrtx E: y=|x|

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

Did you do f(a)+f(b) and f(a+b) for each of the following

OpenStudy (anonymous):

any linear function

OpenStudy (anonymous):

Yes, and I'm not sure which one it would be..

OpenStudy (anonymous):

is it y=2x? :)

OpenStudy (anonymous):

y=x^2 is parabola y=sqrtx is curve y=|x| is like inverted triangle:  so what's left are lines

myininaya (myininaya):

that one is linear and you see if f(a)+f(b)=f(a+b)

myininaya (myininaya):

f(x)=2x f(a)=2a f(b)=2b and f(a+b)=? Then you will see that f(a+b)=f(a)+f(b) after a little distribution

myininaya (myininaya):

But

myininaya (myininaya):

There is also another one

OpenStudy (anonymous):

there is?

myininaya (myininaya):

no only one sorry

OpenStudy (anonymous):

Ok so it is y=2x then?:)

myininaya (myininaya):

yep

myininaya (myininaya):

it has to be linear but you have to watch out to see if has +constant other than 0

OpenStudy (anonymous):

can you help with this? The domain of f(x)=sqrt(x-2)/x^2-x is?

myininaya (myininaya):

\[f(x)=\frac{\sqrt{x-2}}{x^2-x} ?\]

OpenStudy (anonymous):

here are the options.. all reals except 0 and 1 x</= 2, x can't be 0 or 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got all reals, but it can't be 0 or 1.. but is is also x< or equal to 2?

myininaya (myininaya):

Ok the thing under the radical needs to be 0 or positive so that means we need x-2 _?__ 0 What goes where that question mark is... Less than? Greater than? Greater than or equal? Less than or equal? ------- The you also don't want the bottom to be zero

OpenStudy (anonymous):

les than or equal? I'm not sure what you're asking

OpenStudy (anonymous):

its less than or equal as an option

myininaya (myininaya):

No we want x-2 to be zero or positive so that means we want x-2> or = 0

myininaya (myininaya):

\[x-2 \ge 0\] Be we also want: \[ x^2-x \neq 0\]

OpenStudy (anonymous):

So would the answer be x</= 2, then can't be 0 and 1

OpenStudy (anonymous):

I agree with myininaya

myininaya (myininaya):

No no you want x-2 >=0 not <=

OpenStudy (anonymous):

Oh ok so the answer would be x>= 0, 2?? thats it? i thought it can't be 0 and 1

myininaya (myininaya):

You need to solve that inequality I wrote up there

OpenStudy (anonymous):

yeah, its x is greater than or equal to 2.

myininaya (myininaya):

right! :)

OpenStudy (anonymous):

but just that then? or 0 and 2 can't be

myininaya (myininaya):

0 and 2?

OpenStudy (anonymous):

so thats just the answer? because after i factored, i found that 0 and 1 can't be it

OpenStudy (anonymous):

x(x-1)=0, so it can't be 0 and 1..but its just x>=2?

myininaya (myininaya):

you mean the bottom is 0 when x=0 or x=1 right but since x is greater than or equal 2 then we don't have to worry about x=0 or x=1 because x is greater than these numbers

OpenStudy (anonymous):

Ok so you're sure the answer will just be x>/= 2?

myininaya (myininaya):

Am I sure? Am sure if you are sure? Do you need further explanation? Like I mean what is bugging you about this answer?

OpenStudy (anonymous):

I think thats the answer, but is it is what I'm asking? lol

myininaya (myininaya):

Yes if the function is: \[f(x)=\frac{\sqrt{x-2}}{x^2-x}\] Then the domain is \[x \ge 2 \]